find the slope of the line y= 10/3 x-12
Answer by jim_thompson5910(35256)

The slope of y = mx+b is m


Since y = 10/3x-12 is in that form, we can see that m = 10/3


So the slope of the line is 10/3

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim